OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form. Answer y + 3 = 2(x – 4) y – 4 = 2(x + 3) y – 3 = 2(x + 4) y + 4 = 2(x – 3)

OpenStudy (aroub):

y+4=2(x-3)

OpenStudy (anonymous):

its y-4=2(x+3)